The level of thorium in a sample decreases by a factor of one-half every 4.2 million years. A meteorite is discovered to have only 7.6% of its original thorium remaining. How old is the meteorite?
Answer by scott8148(6628)

a formula for half-life is A=Ao(1/2)^(t/h)

A=current amount __ Ao=initial amount __ t=time __ h=half-life

.076=(.5)^(t/4.2x10^6) __ taking log __ log(.076)=(t/4.2x10^6)log(.5)

dividing by log(.5) __ [log(.076)]/[log(.5)]=t/4.2x10^6 __ multiplying by 4.2x10^6 __ {[log(.076)]/[log(.5)]}(4.2x10^6)=t

1.56x10^7=t